P1
Using the average radius of the earth calculate the magnitude of the Earth's gravitational field for a 70 kg person standing on the Earth's surface.

P2
Using the average radius of earth calculate the magnitude earths magnetic field for a 700kg satellite 190km away from the earths surface

P3
Using the earhs average radius calculate the magnitude of the Earth's gravitational field for the moon use [3. 844 x 10^8m for the center to center distance between the earth and the moon
1
Expert's answer
2021-09-06T11:36:56-0400

P1

"F=G\frac{mM_E}{R_E^2}"

"F=6.67*10^{-11}*\frac{70*5.97*10^{24}}{(6378*10^3)^2}=685\: \rm N"

P2

"F=G\frac{mM_E}{R_E^2}"

"F=6.67*10^{-11}*\frac{700*5.97*10^{24}}{((6378+190)*10^3)^2}=6461\: \rm N"

P3

"F=6.67*10^{-11}*\frac{7.35*10^{22}*5.97*10^{24}}{(3. 844 *10^8)^2}=1.98*10^{20}\: \rm N"